A customer wants to make a teepee in his backyard for his children. He plans to use leangths of PVC pipe for the the supports on the teepee, and he wants the teepee to be 12 feet across and 8 feet tall. How long should the pieces of PVC plumbing pipe be?
Answer by checkley77(12844)

WHAT YOU HAVE IS A RIGHT TRIANGLE WITH ONE SIDE (HEIGHT)=8 FT.
THE OTHER SIDE IS 1/2 THE BASE (1/2*12=6 FEET) FOR THE OTHER SIDE.
NOW YOU NEED TO FIND THE HYPOTENUSE OF THIS TRIANGLE USING;
A^2+B^2=C^2
8^2+6^2=C^2
64+36=C^2
100=C^2
C=SQRT100
X=10 FT. FOR THE LENGTH OF THE PIPE FOR THIS TEEPEE.